Given that w denotes either one of the non-real roots of the equation z^3 = 1, show
that
(a) 1 + w + w^2 = 0, and
(b) the other non-real root is w^2
. Show that the non-real roots of the equation

((1 − u)÷u)^3
can be expressed in the form Aw and Bw^2
, where A and B are real
numbers, find A and B.
1
Expert's answer
2020-05-19T18:24:16-0400

"z^3=1\\
z=\sqrt[3]{1}\\
1=\cos0+i\sin0\\
\sqrt[3]{1}=\cos\frac{0+2\pi k}{3}+i\sin\frac{0+2\pi k}{3}, k=0,1,2\\
w_0=\cos0+i\sin0=1\\
w_1=\cos\frac{2\pi}{3}+i\sin\frac{2\pi}{3}=-\frac{1}{2}+i\frac{\sqrt3}{2}\\
w_2=\cos\frac{4\pi}{3}+i\sin\frac{4\pi}{3}=-\frac{1}{2}-i\frac{\sqrt3}{2}\\"

a)

"1+w_1+w^2_1=1-\frac{1}{2}+i\frac{\sqrt3}{2}+\\
+(-\frac{1}{2}+i\frac{\sqrt3}{2})^2=\frac{1}{2}+i\frac{\sqrt3}{2}+\\
+\frac{1}{4}-i\frac{\sqrt3}{2}-\frac{3}{4}=0\\
1+w_2+w^2_2=1-\frac{1}{2}-i\frac{\sqrt3}{2}+\\
+(-\frac{1}{2}-i\frac{\sqrt3}{2})^2=\frac{1}{2}-i\frac{\sqrt3}{2}+\\
+\frac{1}{4}-i\frac{\sqrt3}{2}-\frac{3}{4}=0"

b)

"w_1^2=(-\frac{1}{2}+i\frac{\sqrt3}{2})^2=
\frac{1}{4}-i\frac{\sqrt3}{2}-\frac{3}{4}=\\
=-\frac{1}{2}-i\frac{\sqrt3}{2}=w_2\\
w_2^2=(-\frac{1}{2}-i\frac{\sqrt3}{2})^2=
\frac{1}{4}+i\frac{\sqrt3}{2}-\frac{3}{4}=\\
=-\frac{1}{2}+i\frac{\sqrt3}{2}=w_1"


"(\frac{1-u}{u})^3=1\\
u=Aw_1=A(-\frac{1}{2}+i\frac{\sqrt3}{2})\\
\frac{1-u}{u}=\sqrt[3]{1}\\
\frac{1-u}{u}=w_1\\
\frac{1-Aw_1}{Aw_1}=w_1\\
1-Aw_1=Aw_1^2\\
A(w_1+w_1^2)=1\\
w_1^2=(-\frac{1}{2}+i\frac{\sqrt3}{2})^2=\\
=\frac{1}{4}-i\frac{\sqrt3}{2}-\frac{3}{4}=\\
=-\frac{1}{2}+i\frac{\sqrt3}{2}=w_2\\
A(w_1+w_2)=1\\
A(-1)=1\\
A=-1"

"u=Bw_2=B(-\frac{1}{2}-i\frac{\sqrt3}{2})\\
1-Bw_2=Bw_2^2\\
B(w_2+w_1)=1\\
B=-1"